OpenStudy (anonymous):

sqrt(x + 23) + sqrt(x - 26) = 7

OpenStudy (anonymous):

gotta square twice. ick

OpenStudy (kinggeorge):

So you want to find some x such that \(x+23\) is a perfect square, and \(x-26\) is a perfect square. In other words, you want a perfect square \(n\) such that \(n+23+26=n+49\) is also a perfect square. Fortunately, we can notice that \(n=0\) fulfills our requirements very well. Since 0 is a perfect square, and \(0+49=49=7^2\). Thus, we want \(x-26=0\). Solving this, we get \(x=26\)

OpenStudy (anonymous):

is it because x 26 would be positive

OpenStudy (anonymous):

i am not understanding how the answer is 26

OpenStudy (kinggeorge):

Well, \(x-26\) must be a perfect square, and \(x+23\) must be a perfect square. Now let \(n=x-26\). This means that \(x+23=n+26+23=n+49\). Now we know that \(n\) is a perfect square, and \(n+49\) is a perfect square. From a quick glance, we see that if \(n=0\), both properties hold. From there, we backtrack a little bit to where we know \(n=x-26\). Since \(n=0\), we have that \(x-26=0\) so \(x=26\)

OpenStudy (anonymous):

how do we do this problem: Solve. sqrt(5 x) - sqrt(5 x - 21) = 3

OpenStudy (kinggeorge):

I need to go now. Just post the problem in a new thread, and someone should be able to help you.